# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ Appendix B Tables and Figures
Corn (Grain) CEA 0.673 0.800 0.684 0.726 0.849 0.782 0.738 0.606 0.785 0.753 0.761 0.805 0.771 0.779 0.836 0.760 0.874 0.864 0.861 0.883 0.891 0.811
CWP 0.795 0.744 0.737 0.817 0.766 0.746 0.717 0.748 0.711 0.835 0.748 0.815 0.854 0.890 0.841 0.844 0.807 0.848 0.841 0.827 0.816 0.728
MID 0.587 0.670 0.709 0.687 0.705 0.741 0.733 0.630 0.732 0.692 0.500 0.807 0.769 0.703 0.700 0.730 0.703 0.717 0.730 0.700 0.717 0.669
SEA 0.749 0.794 0.673 0.758 0.746 0.771 0.785 0.656 0.684 0.665 0.707 0.610 0.781 0.800 0.698 0.761 0.922 0.679 0.899 0.910 0.872 0.816
SPL 0.769 0.677 0.647 0.704 0.617 0.681 0.763 0.627 0.686 0.738 0.647 0.710 0.768 0.732 0.805 0.770 0.941 0.750 0.942 0.954 0.949 0.865
SWP 0.858 0.800 0.697 0.824 0.857 0.875 0.880 0.871 0.794 0.720 0.885 0.904 0.838 0.931 0.800 0.877 0.903 0.880 0.926 0.910 0.891 0.838
Wheat (Winter) CEA 0.735 0.790 0.714 0.829 0.705 0.789 0.830 0.737 0.829 0.571 0.729 0.757 0.742 0.751 0.860 0.774 0.909 0.832 0.887 0.898 0.921 0.845
CWP 0.632 0.804 0.667 0.828 0.762 0.766 0.751 0.718 0.802 0.838 0.732 0.892 0.861 0.893 0.878 0.850 0.841 0.847 0.819 0.824 0.732 0.689
MID 0.706 0.771 0.685 0.689 0.681 0.696 0.721 0.662 0.645 0.692 0.664 0.718 0.698 0.690 0.753 0.721 0.690 0.747 0.753 0.721 0.747 0.745
SEA 0.762 0.695 0.598 0.741 0.710 0.677 0.765 0.727 0.651 0.605 0.743 0.714 0.629 0.661 0.722 0.662 0.906 0.790 0.876 0.865 0.850 0.848
SPL 0.734 0.714 0.612 0.697 0.754 0.683 0.763 0.729 0.723 0.713 0.726 0.730 0.687 0.674 0.661 0.712 0.894 0.683 0.895 0.872 0.853 0.807
SWP 0.779 0.760 0.675 0.731 0.766 0.813 0.827 0.776 0.780 0.755 0.742 0.877 0.827 0.811 0.843 0.872 0.813 0.813 0.824 0.865 0.848 0.796
Soybeans CEA 0.685 0.779 0.686 0.733 0.839 0.834 0.815 0.649 0.784 0.727 0.754 0.803 0.822 0.773 0.793 0.720 0.893 0.828 0.883 0.864 0.876 0.810
MID 0.604 0.707 0.712 0.681 0.745 0.769 0.751 0.664 0.792 0.741 0.563 0.857 0.740 0.754 0.745 0.773 0.754 0.734 0.745 0.773 0.734 0.690
SEA 0.721 0.731 0.726 0.702 0.809 0.770 0.780 0.714 0.648 0.663 0.679 0.574 0.800 0.805 0.756 0.770 0.914 0.651 0.890 0.867 0.864 0.705
SPL 0.692 0.676 0.700 0.641 0.702 0.718 0.790 0.635 0.682 0.797 0.654 0.726 0.783 0.713 0.786 0.811 0.885 0.795 0.927 0.885 0.880 0.820
Table B7: MAPE values of XGB models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.295 0.251 0.264 0.295 0.166 0.261 0.189 0.326 0.260 0.324 0.267 0.308 0.175 0.301 0.166 0.300 0.119 0.146 0.122 0.194 0.138 0.262
CWP 0.230 0.269 0.351 0.254 0.279 0.247 0.302 0.318 0.282 0.280 0.369 0.253 0.237 0.236 0.191 0.267 0.234 0.219 0.179 0.237 0.272 0.226
MID 0.282 0.310 0.367 0.312 0.222 0.301 0.273 0.352 0.292 0.314 0.370 0.264 0.250 0.345 0.301 0.312 0.345 0.244 0.301 0.312 0.244 0.387
SEA 0.217 0.298 0.419 0.266 0.251 0.369 0.219 0.231 0.269 0.359 0.326 0.284 0.266 0.309 0.217 0.344 0.086 0.314 0.154 0.145 0.111 0.257
SPL 0.173 0.205 0.282 0.194 0.228 0.270 0.201 0.183 0.215 0.312 0.167 0.154 0.182 0.173 0.225 0.202 0.048 0.208 0.052 0.042 0.057 0.067
SWP 0.112 0.108 0.243 0.038 0.081 0.130 0.088 0.043 0.114 0.274 0.036 0.079 0.190 0.092 0.077 0.140 0.035 0.073 0.052 0.035 0.076 0.017
Wheat (Winter) CEA 0.162 0.237 0.329 0.176 0.226 0.249 0.198 0.228 0.211 0.340 0.277 0.257 0.250 0.233 0.171 0.250 0.140 0.179 0.141 0.199 0.141 0.181
CWP 0.179 0.198 0.177 0.177 0.175 0.182 0.168 0.165 0.245 0.098 0.270 0.110 0.163 0.162 0.128 0.199 0.191 0.157 0.135 0.236 0.222 0.208
MID 0.206 0.301 0.365 0.280 0.271 0.306 0.279 0.372 0.344 0.278 0.369 0.316 0.284 0.337 0.302 0.351 0.337 0.277 0.302 0.351 0.277 0.371
SEA 0.249 0.281 0.379 0.292 0.290 0.317 0.193 0.268 0.251 0.321 0.282 0.261 0.286 0.373 0.127 0.300 0.093 0.244 0.180 0.158 0.147 0.170
SPL 0.167 0.236 0.231 0.213 0.161 0.178 0.207 0.193 0.182 0.239 0.206 0.220 0.210 0.215 0.185 0.185 0.065 0.225 0.060 0.068 0.085 0.101
SWP 0.055 0.102 0.132 0.064 0.075 0.070 0.061 0.092 0.053 0.085 0.071 0.082 0.058 0.104 0.050 0.108 0.087 0.070 0.058 0.091 0.094 0.151
Soybeans CEA 0.185 0.230 0.248 0.269 0.156 0.212 0.177 0.286 0.251 0.275 0.226 0.266 0.140 0.264 0.169 0.229 0.123 0.159 0.092 0.142 0.101 0.231
MID 0.237 0.272 0.314 0.271 0.199 0.271 0.236 0.309 0.261 0.306 0.339 0.236 0.229 0.318 0.273 0.257 0.236 0.236 0.273 0.257 0.318 0.367
SEA 0.222 0.314 0.330 0.249 0.239 0.315 0.235 0.283 0.264 0.304 0.303 0.304 0.258 0.323 0.198 0.328 0.112 0.299 0.117 0.088 0.087 0.265
SPL 0.209 0.225 0.263 0.199 0.210 0.276 0.174 0.214 0.257 0.265 0.191 0.192 0.196 0.191 0.218 0.185 0.068 0.217 0.085 0.071 0.088 0.132
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.206 0.339 0.298 0.307 0.476 0.545 0.416 0.166 0.411 0.335 0.431 0.470 0.497 0.401 0.438 0.427 0.646 0.476 0.619 0.567 0.535 0.316
CWP 0.442 0.343 0.261 0.405 0.291 0.364 0.370 0.387 0.274 0.519 0.409 0.526 0.530 0.557 0.499 0.592 0.421 0.558 0.608 0.556 0.408 0.315
MID 0.151 0.252 0.246 0.255 0.338 0.408 0.388 0.121 0.283 0.318 0.119 0.444 0.249 0.236 0.312 0.268 0.236 0.320 0.312 0.268 0.320 0.205
SEA 0.253 0.239 0.168 0.341 0.285 0.331 0.347 0.113 0.296 0.282 0.158 0.211 0.485 0.379 0.249 0.377 0.658 0.248 0.715 0.671 0.573 0.507
SPL 0.325 0.192 0.192 0.235 0.190 0.236 0.353 0.240 0.337 0.330 0.172 0.336 0.356 0.373 0.402 0.362 0.797 0.366 0.803 0.707 0.825 0.573
SWP 0.582 0.492 0.151 0.571 0.516 0.552 0.622 0.553 0.371 0.250 0.599 0.636 0.617 0.619 0.508 0.549 0.666 0.516 0.641 0.719 0.478 0.520
Wheat (Winter) CEA 0.258 0.489 0.094 0.498 0.335 0.429 0.469 0.354 0.496 0.167 0.308 0.291 0.403 0.354 0.580 0.455 0.618 0.531 0.631 0.709 0.658 0.492
CWP 0.173 0.349 0.201 0.375 0.320 0.285 0.244 0.433 0.349 0.530 0.408 0.629 0.603 0.576 0.547 0.485 0.549 0.533 0.595 0.572 0.392 0.294
MID 0.250 0.302 0.148 0.240 0.190 0.195 0.239 0.206 0.177 0.171 0.172 0.175 0.219 0.175 0.353 0.273 0.175 0.390 0.353 0.273 0.390 0.297
SEA 0.351 0.362 0.135 0.311 0.255 0.258 0.255 0.308 0.278 0.129 0.284 0.236 0.167 0.173 0.323 0.212 0.577 0.216 0.609 0.490 0.488 0.598
SPL 0.284 0.252 0.245 0.241 0.353 0.321 0.298 0.256 0.283 0.341 0.251 0.328 0.363 0.301 0.309 0.335 0.657 0.253 0.666 0.626 0.607 0.518
SWP 0.417 0.365 0.152 0.411 0.345 0.437 0.412 0.408 0.324 0.224 0.308 0.545 0.471 0.380 0.392 0.411 0.570 0.522 0.456 0.560 0.506 0.440
Soybeans CEA 0.226 0.360 0.300 0.323 0.492 0.631 0.494 0.199 0.377 0.311 0.502 0.458 0.516 0.437 0.427 0.435 0.626 0.432 0.630 0.555 0.525 0.370
MID 0.181 0.281 0.276 0.327 0.394 0.478 0.474 0.114 0.386 0.358 0.162 0.504 0.346 0.276 0.320 0.324 0.320 0.371 0.276 0.324 0.371 0.200
SEA 0.155 0.137 0.205 0.247 0.303 0.331 0.332 0.060 0.192 0.287 0.092 0.184 0.465 0.318 0.273 0.383 0.604 0.228 0.534 0.543 0.470 0.302
SPL 0.326 0.229 0.250 0.246 0.239 0.265 0.399 0.209 0.302 0.347 0.183 0.322 0.418 0.385 0.454 0.404 0.684 0.360 0.709 0.617 0.736 0.500
Table B9: MAPE values of LGBM models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.301 0.273 0.325 0.260 0.270 0.255 0.255 0.357 0.222 0.325 0.270 0.309 0.233 0.311 0.247 0.273 0.220 0.287 0.213 0.249 0.243 0.254
CWP 0.246 0.217 0.332 0.267 0.265 0.260 0.264 0.241 0.276 0.277 0.215 0.238 0.207 0.219 0.254 0.312 0.264 0.286 0.197 0.261 0.334 0.281
MID 0.303 0.320 0.322 0.309 0.253 0.293 0.317 0.330 0.301 0.311 0.336 0.280 0.306 0.332 0.323 0.291 0.332 0.275 0.323 0.291 0.275 0.362
SEA 0.327 0.352 0.422 0.355 0.341 0.365 0.320 0.402 0.362 0.367 0.363 0.358 0.283 0.305 0.342 0.379 0.256 0.360 0.242 0.231 0.253 0.267
SPL 0.152 0.216 0.270 0.209 0.172 0.284 0.211 0.168 0.177 0.272 0.194 0.135 0.196 0.172 0.233 0.217 0.080 0.235 0.062 0.101 0.056 0.087
SWP 0.129 0.155 0.242 0.126 0.139 0.195 0.122 0.049 0.129 0.230 0.054 0.096 0.199 0.086 0.086 0.166 0.060 0.153 0.072 0.045 0.073 0.071
Wheat (Winter) CEA 0.250 0.216 0.332 0.189 0.234 0.242 0.213 0.246 0.235 0.319 0.250 0.268 0.230 0.244 0.221 0.278 0.194 0.234 0.184 0.202 0.213 0.196
CWP 0.182 0.215 0.203 0.179 0.201 0.244 0.182 0.187 0.209 0.138 0.237 0.125 0.161 0.132 0.165 0.193 0.173 0.221 0.135 0.239 0.221 0.183
MID 0.257 0.268 0.300 0.262 0.315 0.287 0.268 0.313 0.302 0.303 0.301 0.306 0.322 0.298 0.291 0.267 0.298 0.247 0.291 0.267 0.247 0.344
SEA 0.238 0.262 0.356 0.276 0.325 0.261 0.296 0.260 0.265 0.308 0.301 0.304 0.288 0.284 0.234 0.285 0.166 0.267 0.173 0.252 0.222 0.219
SPL 0.138 0.207 0.237 0.192 0.195 0.228 0.210 0.160 0.210 0.215 0.191 0.186 0.206 0.200 0.165 0.201 0.096 0.198 0.100 0.094 0.098 0.116
SWP 0.081 0.099 0.066 0.064 0.058 0.053 0.040 0.107 0.075 0.063 0.074 0.048 0.064 0.028 0.030 0.083 0.071 0.079 0.048 0.091 0.110 0.105
Soybeans CEA 0.263 0.240 0.295 0.227 0.228 0.209 0.216 0.305 0.204 0.290 0.231 0.279 0.201 0.272 0.225 0.232 0.215 0.251 0.187 0.213 0.196 0.213
MID 0.259 0.289 0.294 0.262 0.214 0.261 0.266 0.308 0.254 0.292 0.299 0.257 0.279 0.293 0.305 0.244 0.233 0.233 0.305 0.244 0.293 0.336
SEA 0.321 0.323 0.331 0.345 0.286 0.295 0.287 0.385 0.345 0.318 0.342 0.321 0.262 0.300 0.307 0.320 0.247 0.297 0.233 0.241 0.248 0.282
SPL 0.152 0.210 0.252 0.218 0.195 0.255 0.211 0.175 0.194 0.238 0.212 0.167 0.170 0.178 0.193 0.190 0.094 0.196 0.115 0.134 0.130 0.132
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.654 0.673 0.673 0.769 0.843 0.860 0.878 0.678 0.712 0.789 0.802 0.867 0.858 0.884 0.950 0.883 0.958 0.946 0.959 0.962 0.960 0.777
CWP 0.749 0.708 0.700 0.826 0.847 0.836 0.859 0.817 0.755 0.850 0.860 0.947 0.918 0.957 0.956 0.962 0.956 0.960 0.958 0.962 0.959 0.730
MID 0.682 0.683 0.750 0.792 0.813 0.826 0.852 0.636 0.743 0.742 0.802 0.841 0.802 0.857 0.927 0.908 0.857 0.931 0.927 0.908 0.931 0.618
SEA 0.746 0.727 0.675 0.801 0.835 0.831 0.878 0.653 0.702 0.634 0.747 0.823 0.836 0.847 0.907 0.866 0.975 0.934 0.979 0.976 0.978 0.863
SPL 0.782 0.736 0.617 0.857 0.843 0.811 0.866 0.694 0.682 0.730 0.762 0.875 0.816 0.883 0.913 0.892 0.983 0.920 0.983 0.983 0.984 0.896
SWP 0.835 0.814 0.560 0.887 0.910 0.885 0.928 0.876 0.774 0.758 0.893 0.925 0.886 0.937 0.946 0.946 0.971 0.955 0.970 0.973 0.974 0.797
Wheat (Winter) CEA 0.711 0.825 0.546 0.864 0.825 0.863 0.893 0.758 0.811 0.711 0.850 0.831 0.869 0.893 0.942 0.908 0.961 0.952 0.970 0.963 0.971 0.846
CWP 0.659 0.746 0.747 0.792 0.827 0.884 0.901 0.801 0.779 0.878 0.851 0.938 0.930 0.955 0.964 0.958 0.959 0.966 0.967 0.960 0.967 0.782
MID 0.687 0.702 0.704 0.802 0.746 0.797 0.848 0.696 0.744 0.698 0.800 0.816 0.819 0.853 0.891 0.876 0.853 0.920 0.891 0.876 0.920 0.625
SEA 0.662 0.757 0.655 0.774 0.789 0.848 0.859 0.611 0.687 0.605 0.769 0.777 0.790 0.818 0.914 0.824 0.962 0.919 0.961 0.960 0.966 0.803
SPL 0.763 0.687 0.651 0.800 0.873 0.847 0.887 0.698 0.725 0.744 0.752 0.837 0.846 0.861 0.890 0.883 0.956 0.905 0.961 0.965 0.965 0.832
SWP 0.739 0.760 0.656 0.852 0.874 0.854 0.897 0.747 0.787 0.807 0.846 0.915 0.898 0.929 0.930 0.944 0.966 0.951 0.970 0.967 0.970 0.748
Soybeans CEA 0.674 0.730 0.718 0.814 0.831 0.890 0.896 0.696 0.724 0.783 0.812 0.871 0.879 0.905 0.942 0.901 0.961 0.942 0.949 0.952 0.961 0.797
MID 0.733 0.720 0.738 0.838 0.818 0.865 0.878 0.649 0.798 0.726 0.799 0.863 0.836 0.884 0.927 0.920 0.927 0.946 0.884 0.920 0.946 0.627
SEA 0.691 0.681 0.744 0.783 0.839 0.829 0.867 0.502 0.717 0.710 0.762 0.772 0.818 0.833 0.922 0.858 0.969 0.938 0.966 0.968 0.971 0.819
SPL 0.752 0.681 0.693 0.841 0.848 0.851 0.902 0.671 0.684 0.781 0.731 0.872 0.850 0.893 0.917 0.895 0.969 0.915 0.964 0.972 0.976 0.845
Table B11: MAPE values of XGB-C models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.247 0.281 0.328 0.280 0.196 0.243 0.206 0.317 0.232 0.373 0.254 0.316 0.249 0.275 0.172 0.285 0.155 0.181 0.150 0.148 0.149 0.237
CWP 0.217 0.222 0.348 0.222 0.257 0.267 0.244 0.250 0.241 0.281 0.225 0.251 0.256 0.235 0.221 0.240 0.234 0.226 0.232 0.245 0.230 0.304
MID 0.271 0.329 0.321 0.302 0.301 0.332 0.329 0.382 0.306 0.373 0.313 0.362 0.334 0.308 0.281 0.310 0.308 0.275 0.281 0.310 0.275 0.398
SEA 0.289 0.269 0.381 0.298 0.293 0.342 0.281 0.347 0.292 0.401 0.291 0.246 0.227 0.192 0.111 0.179 0.051 0.096 0.057 0.056 0.059 0.181
SPL 0.145 0.261 0.283 0.175 0.164 0.218 0.148 0.218 0.232 0.284 0.195 0.154 0.194 0.150 0.224 0.173 0.049 0.225 0.051 0.055 0.055 0.049
SWP 0.027 0.100 0.277 0.048 0.031 0.073 0.042 0.045 0.106 0.246 0.032 0.082 0.088 0.076 0.066 0.077 0.049 0.080 0.043 0.049 0.044 0.050
Wheat (Winter) CEA 0.152 0.172 0.364 0.166 0.170 0.170 0.160 0.268 0.232 0.351 0.244 0.267 0.238 0.236 0.124 0.248 0.156 0.140 0.136 0.145 0.128 0.160
CWP 0.176 0.187 0.201 0.174 0.198 0.186 0.208 0.166 0.162 0.141 0.170 0.152 0.131 0.137 0.113 0.137 0.134 0.120 0.119 0.134 0.117 0.174
MID 0.212 0.317 0.314 0.277 0.274 0.299 0.264 0.375 0.302 0.344 0.269 0.333 0.278 0.270 0.244 0.276 0.270 0.223 0.244 0.276 0.223 0.369
SEA 0.242 0.204 0.350 0.235 0.256 0.236 0.224 0.265 0.238 0.304 0.268 0.203 0.200 0.225 0.121 0.249 0.087 0.112 0.100 0.099 0.101 0.134
SPL 0.143 0.212 0.226 0.180 0.163 0.207 0.173 0.111 0.221 0.251 0.202 0.171 0.225 0.170 0.181 0.178 0.056 0.163 0.063 0.058 0.064 0.088
SWP 0.069 0.088 0.148 0.091 0.078 0.060 0.077 0.067 0.058 0.067 0.094 0.078 0.044 0.086 0.068 0.070 0.087 0.067 0.082 0.088 0.083 0.071
Soybeans CEA 0.199 0.214 0.297 0.217 0.205 0.212 0.187 0.291 0.174 0.331 0.210 0.281 0.225 0.252 0.181 0.251 0.100 0.183 0.100 0.106 0.095 0.199
MID 0.239 0.292 0.297 0.292 0.255 0.302 0.290 0.343 0.299 0.331 0.305 0.321 0.293 0.279 0.275 0.276 0.279 0.249 0.275 0.276 0.249 0.371
SEA 0.326 0.337 0.311 0.392 0.255 0.295 0.280 0.375 0.275 0.308 0.243 0.232 0.239 0.202 0.142 0.195 0.088 0.125 0.091 0.093 0.093 0.228
SPL 0.140 0.259 0.234 0.158 0.172 0.229 0.174 0.166 0.265 0.243 0.185 0.207 0.221 0.201 0.237 0.198 0.068 0.233 0.069 0.086 0.083 0.112

### v2

@@ Appendix B Tables
Train Corn CEA 0.051 0.054 0.047 0.040 0.032 0.034 0.029 0.040 0.055 0.056 0.034 0.030 0.039 0.029 0.029 0.029 0.028 0.029 0.029 0.028 0.028 0.053
CWP 0.045 0.033 0.035 0.031 0.031 0.028 0.028 0.037 0.038 0.027 0.030 0.025 0.024 0.022 0.021 0.022 0.019 0.021 0.019 0.018 0.018 0.032
MID 0.032 0.036 0.038 0.025 0.022 0.028 0.020 0.036 0.034 0.045 0.024 0.027 0.030 0.021 0.020 0.021 0.021 0.020 0.020 0.021 0.020 0.038
SEA 0.039 0.062 0.051 0.035 0.027 0.041 0.025 0.042 0.062 0.049 0.037 0.035 0.039 0.031 0.028 0.028 0.027 0.025 0.025 0.024 0.023 0.057
SPL 0.133 0.134 0.161 0.090 0.105 0.113 0.082 0.160 0.152 0.129 0.134 0.090 0.100 0.080 0.072 0.078 0.033 0.072 0.031 0.032 0.030 0.100
SWP 0.151 0.152 0.191 0.113 0.090 0.098 0.080 0.134 0.154 0.145 0.114 0.082 0.091 0.072 0.067 0.052 0.050 0.051 0.049 0.044 0.044 0.146
Wheat CEA 0.037 0.040 0.036 0.030 0.027 0.026 0.023 0.035 0.039 0.033 0.031 0.024 0.028 0.024 0.023 0.023 0.023 0.022 0.022 0.021 0.020 0.041
CWP 0.066 0.051 0.047 0.046 0.043 0.034 0.033 0.049 0.048 0.034 0.037 0.026 0.025 0.021 0.020 0.021 0.020 0.019 0.019 0.019 0.018 0.043
MID 0.043 0.043 0.038 0.038 0.032 0.033 0.026 0.041 0.049 0.038 0.039 0.031 0.034 0.030 0.028 0.027 0.030 0.025 0.028 0.027 0.025 0.044
SEA 0.042 0.047 0.036 0.036 0.028 0.026 0.024 0.052 0.043 0.045 0.036 0.035 0.036 0.030 0.028 0.023 0.026 0.022 0.025 0.022 0.021 0.050
SPL 0.041 0.047 0.039 0.036 0.030 0.034 0.027 0.040 0.041 0.043 0.030 0.034 0.035 0.027 0.026 0.024 0.028 0.022 0.025 0.025 0.022 0.046
SWP 0.067 0.057 0.056 0.046 0.042 0.037 0.032 0.066 0.055 0.042 0.050 0.033 0.037 0.032 0.027 0.023 0.021 0.021 0.020 0.019 0.018 0.068
Soybeans CEA 0.059 0.052 0.045 0.044 0.036 0.035 0.033 0.050 0.058 0.057 0.042 0.040 0.045 0.035 0.033 0.031 0.032 0.031 0.032 0.028 0.029 0.060
MID 0.037 0.034 0.029 0.030 0.024 0.024 0.023 0.031 0.026 0.031 0.024 0.023 0.024 0.020 0.017 0.020 0.020 0.017 0.017 0.020 0.017 0.030
SEA 0.040 0.059 0.049 0.035 0.029 0.033 0.026 0.041 0.058 0.054 0.029 0.034 0.036 0.025 0.024 0.024 0.020 0.022 0.019 0.019 0.019 0.054
SPL 0.036 0.043 0.035 0.027 0.025 0.026 0.023 0.044 0.044 0.046 0.032 0.037 0.034 0.027 0.022 0.027 0.025 0.023 0.022 0.026 0.023 0.054
Test Corn CEA 0.109 0.119 0.087 0.101 0.093 0.108 0.093 0.107 0.108 0.104 0.102 0.096 0.113 0.103 0.104 0.104 0.101 0.106 0.099 0.102 0.099 0.109
CWP 0.044 0.046 0.040 0.047 0.043 0.039 0.041 0.037 0.040 0.043 0.041 0.045 0.043 0.044 0.045 0.045 0.046 0.048 0.045 0.044 0.044 0.049
MID 0.047 0.059 0.045 0.048 0.042 0.051 0.045 0.043 0.049 0.049 0.044 0.042 0.043 0.042 0.046 0.042 0.046 0.046 0.046 0.042 0.042 0.058
SEA 0.068 0.089 0.075 0.070 0.066 0.074 0.065 0.074 0.088 0.080 0.073 0.069 0.071 0.069 0.067 0.072 0.073 0.071 0.073 0.072 0.072 0.086
SPL 0.210 0.315 0.317 0.233 0.223 0.294 0.236 0.254 0.281 0.275 0.272 0.239 0.221 0.228 0.246 0.237 0.102 0.249 0.089 0.095 0.084 0.107
SWP 0.137 0.176 0.332 0.141 0.151 0.187 0.154 0.095 0.192 0.291 0.104 0.140 0.189 0.144 0.150 0.190 0.157 0.187 0.115 0.157 0.111 0.169
Wheat CEA 0.066 0.067 0.062 0.069 0.063 0.057 0.059 0.068 0.063 0.069 0.066 0.067 0.060 0.063 0.066 0.064 0.066 0.063 0.063 0.068 0.060 0.066
CWP 0.102 0.087 0.094 0.093 0.086 0.081 0.086 0.082 0.092 0.091 0.080 0.078 0.084 0.079 0.079 0.079 0.074 0.076 0.079 0.076 0.079 0.091
MID 0.068 0.075 0.069 0.070 0.071 0.065 0.067 0.082 0.078 0.069 0.083 0.068 0.069 0.069 0.071 0.070 0.069 0.071 0.071 0.070 0.071 0.078
SEA 0.083 0.078 0.060 0.079 0.069 0.072 0.073 0.078 0.088 0.066 0.079 0.069 0.079 0.074 0.080 0.072 0.084 0.079 0.071 0.083 0.070 0.075
SPL 0.087 0.091 0.074 0.081 0.080 0.073 0.075 0.083 0.079 0.071 0.066 0.077 0.080 0.071 0.072 0.070 0.072 0.071 0.067 0.075 0.067 0.100
SWP 0.143 0.148 0.154 0.149 0.150 0.154 0.149 0.140 0.128 0.146 0.134 0.146 0.152 0.148 0.153 0.160 0.145 0.146 0.153 0.154 0.146 0.157
Soybeans CEA 0.106 0.115 0.077 0.100 0.096 0.083 0.090 0.097 0.112 0.094 0.106 0.096 0.093 0.098 0.093 0.108 0.080 0.102 0.090 0.098 0.077 0.112
MID 0.054 0.066 0.046 0.054 0.054 0.061 0.058 0.061 0.060 0.055 0.061 0.056 0.054 0.058 0.056 0.058 0.047 0.054 0.056 0.058 0.045 0.059
SEA 0.105 0.127 0.100 0.109 0.076 0.086 0.074 0.103 0.129 0.105 0.100 0.095 0.106 0.091 0.092 0.094 0.088 0.093 0.074 0.076 0.087 0.120
SPL 0.073 0.090 0.066 0.082 0.075 0.063 0.086 0.085 0.096 0.073 0.088 0.083 0.087 0.081 0.077 0.076 0.058 0.069 0.077 0.080 0.064 0.085